# Webly-Supervised Image Manipulation Localization via Category-Aware Auto-Annotation

### v1

@@ III-A Self-Supervised Classification
III-B Difference-Aware Semantic Segmentation (DASS)
@@
Ideally, for an SPG image pair, the absolute difference between the authentic and forged images would directly reveal the forged region. However, due to transmission noise and image degradation [28], the absolute difference map is often imprecise. As shown in Fig. 4, almost every area in the difference map shows nonzero values because of noise. Even after binarizing the difference map using an OTSU [29] adaptive threshold, noisy highlights persist, especially in high-frequency regions such as edges. To address this issue, we propose denoising the difference map by incorporating semantic information from the images. Specifically, we concatenate the authentic image, the forged image, and their absolute difference map along the channel dimension, and feed the result into a semantic segmentation model. Our segmentation model employs a VAN [30] encoder for semantic feature extraction and utilizes the same Multi-Aspect Denoiser as in our Corr-DINO to denoise the difference map.
III-C Correlation DINO (Corr-DINO)
CIML on SDG requires detecting shared objects at various scales within complex backgrounds. This task is considerably more challenging than SPG, where targets are already highlighted through image difference maps. As a result, synthetic or small-scale handcrafted datasets are largely insufficient for training a robust model for complex real-world scenarios.
To tackle the overfitting issue, we introduce a novel frozen-denoising paradigm (see Fig. 5). Our key insight is that overfitting can be mitigated by employing a frozen encoder and subsequently denoising its output features. Building on this paradigm, we developed the Corr-DINO model (see Fig. 6), which consists of five components:
1. Frozen Encoder for feature extraction. We extract features from both the tampered image $I_{a}$ and its original image $I_{b}$ with a frozen DINOv2 [13] backbone. This ViT backbone has learned generalized representations through self-supervised learning on vast-scale web images. The features from the last ViT layer, denoted as $F_{a1}$ and $F_{b1}$, contain rich semantic information and are used to compute correlation features.
2. Correlation module for correlation features calculation. $F_{a\_corr}=[Corr(F_{a1},F_{b1}),Corr(F_{a1},F_{a1})]$, $F_{b\_corr}=[Corr(F_{b1},F_{a1}),Corr(F_{b1},F_{b1})]$. The same correlation function $Corr$ as most previous works [12, 27, 14] is adopted. The detailed $Corr$ function is presented in the Appendix.
Figure 5: Different from previous works that employ learnable encoders and rule-based feature aggregation for constrained image manipulation localization on SDG image pairs, we propose a novel frozen-denoising paradigm. This paradigm offers significantly better generalization in real-world scenarios.
Figure 6: The overall pipeline of our proposed Correlation DINO. First, we employ a frozen DINO backbone to extract image features. The features obtained from the final DINO layer are then used for correlation calculation. Next, a learnable aggregation module reduces the channel dimension of the correlation features, which are further refined using a Feature Super Resolution module. Finally, the refined features are denoised with a Multi-Aspect Denoiser, resulting in the final mask prediction. Green lines indicate operations used only during training with synthetic data, due to the lack of mask labels for the real images.
3. A Learnable Aggregation module for dimension reduction. Previous works simply select the top-K activated channels with the highest average values in $F_{a\_corr}$ and $F_{b\_corr}$ to reduce dimension. This approach works for trainable encoders but is suboptimal for our frozen encoder, whose correlated object features are relatively sparse. Selecting only the top-K channels may discard important features and compromise the integrity of $F_{a\_corr}$ and $F_{b\_corr}$. To overcome this, we propose a learnable aggregation module that uses trainable parameters to reduce the channel count to K. Following prior methods, we also include statistics such as average and maximum values across the channel dimension. $F_{a\_aggr}=[C_{1\times 1}(R(C_{1\times 1}(F_{a\_corr}))),Avg(F_{a\_corr}),Max(F_{a\_corr})]$. Here, $C_{1\times 1}$ is $1\times 1$ conv-layer and $R$ is ReLU. A similar process is used for $F_{b\_corr}$ to generate $F_{b\_aggr}$.
4. Feature Super Resolution module for image detail reconstruction. The features from the last ViT layer are rich in semantics but lack fine image details [13], making $F_{a\_aggr}$ and $F_{b\_aggr}$ too coarse for fine-grained mask predictions. To reconstruct image details, we leverage features from the last 2-4 layers of DINO, which are known to be more sensitive to edges and fine details [13]. We concatenate the features from the last four encoder layers $F_{a}=Concat(F_{a1,2,3,4})$ and interpolate $F_{a\_aggr}$ to different scales ($1\times,2\times,4\times,8\times$) to obtain $F_{aggr\_a1,2,3,4}$. We then fuse them with $1\times 1$ conv-layers. $F_{asr\_i}=Conv_{1\times 1}(Conv_{1\times 1}(F_{a})+Conv_{1\times 1}(F_{aggr\_ai}))$ for iin(1,2,3,4)i\,in\,(1,2,3,4). A similar process is applied to $F_{b\_aggr}$.
5. Multi-Aspect Denoiser module for features denoising. Since the ViT encoder is frozen, its output features inevitably contain significant task-irrelevant noise. To achieve precise mask prediction, we propose denoising the correlation features by integrating multi-perspective information. This module fuses multi-scale features $F_{asr\_1,2,3,4}$ in a top-down manner. $F^{\prime}4=Conv_{3\times 3}([F_{asr\_4},Avg(F_{asr\_4})])$, $F^{\prime}_{n}=Conv_{3\times 3}(F^{\prime}_{n+1}+Conv_{1\times 1}(F_{asr\_n}))$ for iin(1,2,3)i\,in\,(1,2,3). Here, $Conv_{3\times 3}$ and $Conv_{1\times 1}$ are conv-layers with kernel size 3 and 1, and $Avg$ is global average pooling. The fused features are further concatenated and denoised using a set of dilated $3\times 3$} conv-layers. The output is concatenated to $F^{\prime}1$ and processed by a $3\times 3$} conv-layer to yield the mask prediction.
Our Correlation DINO is optimized end-to-end using cross-entropy loss. It significantly alleviates the overfitting problem in challenging real-world CIML tasks for SDG.
@@ IV MIMLv2 Dataset
In this section, we introduce MIMLv2, a large-scale, diverse, and high-quality dataset. The key idea is to leverage constrained image manipulation localization models trained on existing datasets to automatically annotate manually forged images collected from the web. To ensure high annotation quality, we propose a novel metric that filters out inadequate annotations.
@@
Figure 7: The overall construction pipeline of our MIMLv2 dataset.
@@ IV-A Dataset Construction
Image Collection. We collect forged images and their corresponding originals from imgur.com. On this website, the images are manually manipulated and thus have high-quality, diverse forged regions. Additionally, we exclude images that already appear in the evaluation datasets (Section VII) and remove duplicates by checking their MD5 and pHash.
Auto-Annotation. We utilize the CAAAv2 models, as proposed in Section III, to obtain mask annotations for the collected images. These models are sufficiently trained with one million synthetic COCO samples [26], along with the handcrafted datasets CASIAv2 [31] and IMD20 [32].
@@
Figure 8: Some images and their corresponding mask annotations from the proposed MIMLv2 dataset. Left: SPG image pairs. Right: SDG image pairs.
@@ IV-B Dataset Highlights
We present a few samples of the proposed MIMLv2 dataset in Fig. 8. The main highlights of our dataset are as follows:
  • High quality. Image manipulations in the proposed dataset are meticulously crafted by humans. Such data can teach models to detect real-world forgery, rather than merely overfitting to simple patterns found in synthetic data.
  • Large Scale. As shown in Table I, the proposed dataset comprises 246,212 manually forged images. This is dozens of times more samples than previous handcrafted IML datasets (e.g., $\approx$120 times more samples than IMD20 [32]).
  • Board Diversity. MIMLv2 comprises images of various sizes, styles, and manipulation types (e.g., copy-move, splicing, removal, AIGC). These images are created by hundreds of thousands of individuals using various software. Such broad diversity can considerably enhance the generalization ability of deep IML models.
  • Modern Style. MIMLv2 contains abundant modern images that were recently captured and forged, reflecting current photography and forgery technologies. In contrast, the CASIA dataset [31] proposed over a decade ago, features blurred small images and outdated forgeries. Therefore, MIMLv2 better meets the current requirements for IML.
@@ V Object Jitter
We experimentally find that the IML model performance can be further improved with more high-quality data. To this end, we introduce Object Jitter, an effective method that enhances IML models by generating high-quality image artifacts online during training from authentic web images.
Previous works synthesize tampered images by randomly cropping and pasting image objects [37, 9]. However, these data synthesis methods often produce overly obvious artifacts and semantically unreasonable content, as shown on the left in Fig. 9. Models trained on such data typically overfit to simple cases and perform poorly on real-world forgeries, where artifacts are subtle and tampered content is semantically reasonable. DEFACTO [8] attempts to ensure semantic reasonableness by designing elaborate rules. It manually identifies eight object types with consistently similar shapes and appearances (e.g., the road sign in the middle of Fig. 9); copy-move or splicing operations are restricted to objects of the same type. Because DEFACTO applies only to a limited number of object classes, the generated data lacks diversity, causing models to overfit the semantics of these specific object types. Furthermore, DEFACTO heavily relies on accurate object class annotations, which significantly limits its broader application.
To overcome these challenges, we propose a novel idea: instead of altering image semantics through tampering, we simulate subtle artifacts. Building on this idea, we introduce Object Jitter, a novel method detailed in Fig 10. Given an input image, we readily obtain a segmentation mask for a random object using off-the-shelf segmentation models like SAM [38]. The selected object is then processed using one or a combination of the following three random operations:
  • Size Jitter. Slightly enlarge the selected object’s region and its corresponding mask by a random scale.
  • Exposure Jitter. Slightly overexpose the selected object.
  • Texture Jitter. Slightly alter the texture of the selected object region by applying a random combination of JPEG compression, reverse JPEG compression [39], and blur.
After processing, we add a random smooth transition on the edges to minimize visual artifacts. We apply Object Jitter to 1 to 3 random objects within an input image. The mask regions of these processed objects are recorded and subsequently considered as tampered regions during training.
Our method successfully addresses existing issues: First, it always preserves the semantics of the input image, thereby preventing models from overfitting to simple cases with abnormal semantics. Second, by applying subtle changes to objects that already naturally blend into the background, the generated visual image artifacts are unobvious. Third, our method is unconstrained and universally applicable. Consequently, Object Jitter consistently produces high-quality artifacts and effectively supplements our MIMLv2 dataset.
Figure 9: Comparison between different data generation methods. Our method is universally applicable and always produces semantically reasonable forgeries.
Figure 10: The proposed Object Jitter method. Given a randomly selected object from an authentic image, we slightly enlarge or overexpose it, or alter its texture. The processed object is regarded as tampered object for training.
@@ VI Web-IML
Figure 11: The overall framework of the proposed Web-IML model.
@@
Multi-Scale Perception module motivation. During meticulous image forensic analysis, humans often zoom in and out the image repeatedly, comparing between diverse observations to assist their final prediction. To mimic the human perception way, we introduce the Multi-Scale Perception module.
@@
Multi-Scale Perception module pipeline. Specifically, given four feature maps extracted from the backbone model $F_{1},F_{2},F_{3},F_{4}$}, we first extract global features in different scales: $G_{i}=Conv_{1\times 1}(Avg_{i}(F_{4}))\,for\,i\,in\,(1,2,3,6)$, where $Conv_{1\times 1}$ denotes conv-layer with kernel size 1 and $Avg_{i}$ denotes average pooling function with output size $i\times i$. The global features are then resized, concatenated to $F_{4}$ and fused, dimension reduced to get $F^{\prime}4$. $F^{\prime}4=Conv_{3\times 3}([G_{1},G_{2},G_{3},G_{6},F_{4}])$. Next, the high-level features are fused with the low-level features recursively. $F^{\prime}_{n}=Conv_{3\times 3}(F^{\prime}_{n+1}+Conv_{1\times 1}(F_{n}))\,for\,i\,in\,(1,2,3)$. This process fully integrates high-level semantics with low-level image details, resulting in the multi-scale features $F^{\prime}_{1},F^{\prime}_{2},F^{\prime}_{3},F^{\prime}_{4}$}.
Self-Rectification module motivation. Due to the challenging nature of the image manipulation localization task, forensic models inevitably overlook key information and make mistakes. To minimize negligence, we propose the Self-Rectification module. With this module, our model can learn to check and correct its initial predictions.
@@
Self-Rectification module pipeline. As shown at the bottom of Fig. 11, we obtain the initial mask prediction $P_{n}$ from $F^{\prime}_{1}$ via a $\times1$ conv-layer. We then adaptively obtain a rectification feature $F_{r}$ through a tiny Rectification Feature Mapping module ($RFM$), which consists of 5 cascaded conv-layers. $F_{r}=RFM(P_{n})$. The $F_{r}$ informs the model of suspected tampered regions and potentially incorrect prediction regions. Then $F_{r}$ is fused with $F_{a}$ with a $\times3$ conv-layer to get $F_{c}$. $F_{c}=Conv_{3\times 3}([F_{a},P_{c}])$. To enable in-depth, meticulously analyses of $F_{c}$, we introduce the Nested Channel Attention module. As shown in the top left of Fig. 11, it can be formulated as $G_{1}=Avg(F_{c})W_{1}$, $G_{2}=G_{1}W_{2}$, $A_{2}=\sigma(W_{3}G_{2})$, $A_{1}=\sigma(W_{4}G_{1}*A_{2})$, $F_{o}=F_{c}*A_{1}$. Here, $G_{1,2}$ are global vectors, $A_{1,2}$ are attention maps, $W_{1,2,3,4}$ are learnable parameters, $Avg$ is global average pooling and $\sigma$ is the sigmoid function. The output $F_{o}$ is further processed by four $\times3$ conv-layers with dilation sizes $1,2,3,6$ respectively, the output features are concatenated and fused with a $\times3$ conv-layers. The mask prediction $P_{f}$ is obtained from $F_{o}$ with a $\times1$ conv-layer. To get the refined final prediction $P_{f}$, we replace $P_{n}$ with $P_{f}$, replace $F_{a}$ with $F_{o}$, then repeat the self rectification once again.
Our Web-IML is optimized with the cross entropy loss calculated between the prediction mask and the ground-truth.
@@
TABLE II: Comparison on image manipulation localization. ”WS” is our proposed web supervision, which trains models using data from the MIMLv2 and Object Jitter. ”Avg*” is the average score excluding IMD20. ”Param.” is the number of model parameters.
@@
TABLE V: Ablation study on our Web Supervision method with Web-IML model. ‘PSCC-Synthetic‘ denotes the synthetic data from the PSCC-Net [37]. ‘O.J.‘ denotes the proposed Object Jitter method. ‘C.G.‘ denotes the CocoGlide [10]. $\Delta$‘ denotes improvement.
@@
TABLE VI: Ablation study on our Object Jitter method. ‘COCO‘ denotes using the COCO [49] dataset images for Object Jitter. ’SA’ denotes using 400k random images from the SA-1B [38] dataset for Object Jitter. ‘C.G.‘ denotes the CocoGlide dataset [10].
Num Object Jitter Operation
Image
Source
CASIAv1 [31] Coverage [34] NIST16 [35] IMD20 [32] C.G. [10] CIMD [40] MIFD [41] Average
@@
VII Experiments
VII-A Experiments on the Web-IML Model (IML Task)
Implementation Details. For Web-IML, we adopt ConvNeXt-Base [50] as the feature extractor. The model is trained for 200k iterations with a batch size of 16, using a $\times512$ input size, consistent with previous works [7, 10]. We employ Cross-Entropy loss and AdamW [14] optimizer, with the learning rate linearly decaying from 1e-4 to 1e-6. CASIAv2 [31] and tampCOCO [7] serve as baseline training sets. For fair comparison, mask predictions are binarized with a fixed threshold of 0.5 and evaluated using IoU and binary F1-score.
Comparison Study. Table II shows that our Web-IML significantly outperforms existing methods on multiple real-world image forgery benchmarks. For example, it achieves an average IoU that is 24.1 points higher than TruFor. Furthermore, Web-IML consistently outperforms our conference version, APSC-Net, across various settings. These results demonstrate the strong generalization capabilities of our Web-IML model. Qualitative comparison is illustrated in Fig. 12.
Ablation Study. The ablation results are shown in Table III. Setting (2), which includes the proposed Multi-Scale Perception (MP) module, achieves 5.2 points higher IoU than the baseline (1). This indicates that the proposed MP module effectively integrates high-level and low-level features to detect diverse image artifacts. Setting (3), which further incorporates the Self-Rectification (SR) module, yields an additional 3.9 point IoU improvement over Setting (2), as SR enables the model to detect and correct its errors. Setting (5) outperforms (4) via our nested attention module that enables in-depth analysis. Setting (6) outperforms (5) by including a second round of self-rectification refinement, demonstrating that further refinement can minimize incorrect predictions.
VII-B Experiments on the Web-Supervision Method (IML Task)
This subsection evaluates the effectiveness of our proposed web-supervision approach in addressing the data scarcity issue for IML. Our web-supervision encompasses both the MIMLv2 dataset and the Object Jitter method. We compare the performance of the same IML model trained on data created by our web-supervision approach against other synthetic methods.
Implementation Details. The implementation details remain the same as in the previous subsection VII.A. When supplementing the baseline training sets with additional IML training sets, the total training volume is kept constant with a uniform sampling ratio across all datasets.
Comparison Study for the MIMLv2 Dataset. Table V presents the IML performance of the Web-IML model when trained on different datasets. The baseline (1) uses the two commonest training sets: tampCOCO [7] and CASIAv2 [31]. Incorporating synthetic training data from PSCC-Net [37], DEFACTO [8] or OpenSDID [48] diminishes the average model performance (settings 2, 3, 4). This is because the first two synthetic datasets do not notably surpass the baseline data in quality or diversity, while the AIGC forgeries from OpenSDID differ significantly from non-AIGC forgeries. These results confirm that merely increasing training data scale does not lead to improved performance. In contrast, including the MIMLv1 (setting 5) or MIMLv2 (setting 6) datasets significantly improves model performance by an average of 10 and 12 points, respectively, without increasing training or inference burden. These results confirm that our MIMLv1 and MIMLv2 datasets considerably alleviate data scarcity and overfitting issues in real-world scenarios. This is attributable to their large scale, manual forgery, and considerably higher quality and diversity compared to other datasets. Notably, MIMLv2 yields greater improvement due to its larger quantity of high-quality samples and superior diversity compared to MIMLv1.
Comparison Study for the Object Jitter. Table V shows that including the training data generated by our Object Jitter method further improves the model when combined with MIMLv1 (setting 7 outperforms 5) or MIMLv2 (setting 9 outperforms 6) datasets, effectively supplementing them. This is attributed to our Object Jitter method’s consistent generation of diverse, high-quality samples, which further enriches the training data.
Ablation Study for the Object Jitter. Table VI presents the ablation study for Object Jitter operations and image source. Setting (1) is the baseline, where Web-IML is trained with MIMLv2 but without Object Jitter. Setting (7) represents the full Object Jitter method. Settings (2, 3, 4) involve removing the size jitter, texture jitter, and exposure jitter operations, respectively; all of them result in worse performance than setting (7). These results demonstrate that all three operations contribute to improved performance through increased data diversity. Notably, Setting (2), which lacks size jitter, is significantly worse than (7). Meanwhile, Setting (5), which uses only size jitter, performs close to (7). This indicates that the size jitter operation is the primary contributor to the improvement. Since most Internet forgeries are not generated by copy-move, they exhibit exposure and texture artifacts. The size jitter operation, without altering the object’s exposure or texture, reduces model’s reliance on exposure or texture cues, thus better supplementing MIMLv2. Furthermore, Setting (6) applies the full Object Jitter method using COCO [49] as the image source, instead of 400k random images from SA [38] in setting (7). The similar performance of Settings (6) and (7) demonstrates that the Object Jitter method is robust to the choice of image source.
Conclusions Verification on Other IML Models. We extended our verification process to other IML models. Specifically, we evaluated MIMLv2 and Object Jitter on PSCC-Net [37], CAT-Net [7], and TruFor [10] using their official model code. Table VII presents the results, which clearly indicate that MIMLv1, MIMLv2, and Object Jitter all lead to significantly higher scores. MIMLv2, in particular, consistently yields significantly greater improvement than MIMLv1 (conference version). Moreover, Object Jitter further boosts performance when integrated with MIMLv2. These results robustly confirm our previous analyses and conclusions.
@@ VII-B Experiments on the Web-Supervision Method (IML Task)
TABLE VII: Comparison study of our web-supervision on IML task. Blue number for improvement and Red number for decrease. ’B.L.’ denotes baseline setting using tampCOCO [7] and CASIAv2 [31] datasets as training data. ’PSCC-Synth.’ denotes the synthetic dataset of PSCC-Net [37], ’AIGC-Synth.’ denotes the AIGC forgery dataset OpenSDID [48]. ’O.J.’ denotes our Object Jitter method.

### v2

@@ III-C Corr-DINO
III-D Difference-Aware Semantic Segmentation (DASS)
SPG pairs are characterized by in-place manipulation, where the manipulated image is generated by directly altering its original. This process inherently ensures precise spatial alignment and a shared background, meaning that the forged regions are always highlighted by the image difference. Leveraging this strong correspondence, we use the difference map as a prior feature to guide model learning. Given a manipulated image $I_{m}$ and its original $I_{o}$, the absolute difference map $D_{mo}=abs(I_{m}-I_{o})$ can roughly highlight forged regions.
@@ III-D Difference-Aware Semantic Segmentation (DASS)
However, transmission noise and image degradation often contaminate the difference map, making it too noisy for precise localization. As illustrated in Fig. 4, nearly every area exhibits nonzero values due to noise, and even adaptive binarization [43] fails to denoise. To address this issue, we propose denoising the difference map by incorporating semantic information from the images. Specifically, we concatenate $I_{m}$, $I_{o}$, and $D_{mo}$ along the channel dimension, and feed them into a semantic segmentation model. This model integrates a VAN [44] encoder for extracting rich semantic features and employs the same Multi-Aspect Denoiser (III-C) as in our Corr-DINO for denoising the difference map.
@@ IV MIMLv2 Dataset
In this section, we introduce MIMLv2, a large-scale, diverse, and high-quality dataset. The key idea is to leverage CIML models trained on existing datasets to auto-annotate manual forgeries collected from the web, thereby addressing the data scarcity in IML. To ensure high annotation quality, we propose a novel metric that filters out inadequate annotations without requiring ground-truth for evaluation.
@@
Figure 5: The construction pipeline of our MIMLv2 dataset.
@@ IV-A Dataset Construction
1. Image Collection. We collect forged images and their corresponding originals from imgur.com. On this website, the images are manually manipulated and thus have high-quality, diverse forged regions. Additionally, we exclude damaged, duplicated images and those that already appear in the evaluation datasets [45, 46, 47, 48, 49, 50, 51] with MD5 [52] and pHash [53].
2. Auto-Annotation. We utilize the CAAAv2 models, as proposed in Section III, to obtain mask annotations for the collected images. The models are effectively trained on synthetic COCO [40], CASIAv2 [46] and IMD20 [45] datasets.
@@
Figure 6: Some images and their corresponding mask annotations from the proposed MIMLv2 dataset.
@@ IV-B Dataset Highlights
We present a few samples of our MIMLv2 dataset in Fig. 6. The forged regions are precisely annotated with high-quality masks, demonstrating the effectiveness of our auto-annotation approach. The main highlights of our dataset include:
  • High quality. The proposed dataset contains high-quality image manipulations crafted by humans. Such data is crucial to training models that can detect real-world forgeries rather than overfit to synthetic data.
  • Large Scale. The proposed dataset is substantially larger than existing handcrafted IML datasets. As shown in Table I, it comprises 246,212 manually forged images, which is about 120 times more than IMD20 [45].
  • Broad Diversity. MIMLv2 comprises images of various sizes, styles, and manipulation types (e.g., copy-move, splicing, removal, AIGC), created by more than 10k individuals and various software. This diversity significantly enhances the generalization of IML models.
  • Updated Forgeries. MIMLv2 contains numerous modern images that were recently captured and forged, reflecting current photography and forgery technologies. In contrast, the CASIA dataset [46] proposed over a decade ago, features blurred small images and outdated forgeries. Therefore, our MIMLv2 better suits current IML requirements.
@@ V Object Jitter
Sophisticated copy-move forgeries, where an object is copied and pasted within the same image, are difficult to detect. They introduce only subtle edge artifacts while preserving source texture and noise patterns. Their scarcity in web-sourced data leads models to rely more on common noise and compression anomalies, causing them to fail on these hard cases. Existing synthesis methods, designed to generate training data for such forgeries, are flawed: random copy-paste techniques [56, 34] often create semantically incoherent artifacts (e.g., a train on food, Fig. 7), while more constrained approaches like DEFACTO limit tampering to a few objects of consistent shape, lacking object diversity. Both issues can cause models to overfit to unrealistic cues.
To address this gap, we propose Object Jitter, a novel strategy that simulates these forgeries by subtly distorting existing objects. The core operation is Size Jitter: we slightly enlarge a randomly selected object, covering its original footprint. This mimics a realistic copy-move forgery by preserving semantics and texture while introducing only faint boundary artifacts. As illustrated in Fig. 9, the process involves (1) segmenting an object using ground truth or SAM [57], (2) scaling the object and its mask by a small random factor, and (3) applying a smooth alpha-blended transition at the new boundary to minimize extraneous visual cues.
To further enhance data diversity, we introduce two additional minor perturbations: Exposure Jitter (slight overexposure) and Texture Jitter (subtle modifications via JPEG compression, reverse JPEG [58], and blur operations). By preserving semantic plausibility and being applicable to any object, Object Jitter provides a source of challenging training examples. This compels the model to learn the subtle boundary artifacts characteristic of sophisticated forgeries, thereby improving its robustness and generalization.
Figure 7: Comparison between different data generation methods.
Figure 8: The proposed Object Jitter method. Given a randomly selected object from an authentic image, we slightly enlarge or overexpose it, or alter its texture. The processed object is regarded as a tampered object for training.
@@ VI Web-IML
Figure 9: The overall framework of the proposed Web-IML.
VI-A Multi-Scale Perception Module
@@
Motivation. This module is designed to mimic the meticulous image forensic analysis, where humans often zoom in and out the image repeatedly, comparing diverse observations to assist their final prediction.
@@
Pipeline. Given the feature maps $F_{1},F_{2},F_{3},F_{4}$} extracted from the encoder, we first extract global features of different scales through multi-scale average pooling. The global features are then fused and dimension-reduced to get $F^{\prime}4$. Next, the high-level features are fused with the low-level features recursively. This process fully integrates high-level semantics with low-level image details, resulting in the multi-scale features $F^{\prime}_{1},F^{\prime}_{2},F^{\prime}_{3},F^{\prime}_{4}$}. More details are provided in the Appendix.
VI-B Self-Rectification Module
Motivation. This module is designed to make the model learn to check and correct its initial predictions, thereby minimizing the overlooking of key information and mistakes. The Self-Rectification (SR) module improves upon the Self-Calibration (SC) module in our conference paper from two aspects: First, in SC, we map the initial prediction mask to a convolution kernel, use it to convolve the mask and multiply the result by the multi-scale features. Despite its effectiveness, learning the map function to a convolution kernel is difficult, and directly multiplying the convolved mask often causes information loss in the black mask region. In our SR, we map the initial prediction mask to a feature map, and concatenate it with the multi-scale features, addressing the issues. Second, we introduce a Nested Channel Attention module that enables more meticulous analyses of the rectification feature.
@@ VI-B Self-Rectification Module
Pipeline. As shown at the bottom of Fig. 9, we obtain the initial mask prediction $P_{n}$ from $F^{\prime}_{1}$. We then adaptively obtain a rectification feature $F_{r}$ through a tiny Rectification Feature Mapping module ($RFM$), which consists of 5 cascaded conv-layers. The $F_{r}$ informs the model of suspected tampered regions and potentially incorrect prediction regions. $F_{r}$ is fused with $F_{a}$ to get $F_{c}$. Then, in the Nested Channel Attention module, we obtain the global vectors $G_{1}$ from $F_{c}$, and subsequently higher-level global vectors $G_{2}$ from $G_{1}$. Then $G_{1}$ is weighted by the attention map calculated by $G_{2}$ and the results are further used to calculate the channel attention weights for $F_{c}$. The output $F_{o}$ is further processed by four dilated conv-layers, the output features are concatenated and fused to produce the mask prediction $P_{f}$. To get the refined final prediction $P_{f}$, we replace $P_{n}$ with $P_{f}$, replace $F_{a}$ with $F_{o}$, then repeat the self-rectification once more.
Our Web-IML is optimized with cross entropy loss in an end-to-end manner.
@@
TABLE II: Comparison on image manipulation localization. ‘WS‘ is our proposed web supervision, which trains models using data from the MIMLv2 and Object Jitter. ‘Avg*‘ is the average score excluding IMD20. ‘Param.‘ is the number of model parameters. CAT-Net, TruFor and SparseViT were trained with the IMD20 dataset and therefore not tested on it.
@@
TABLE V: Ablation study on our Web Supervision method. ‘PSCC-Synthetic‘ denotes the synthetic data from the PSCC-Net [56]. ‘O.J.‘ denotes the proposed Object Jitter method. ‘C.G.‘ denotes the CocoGlide [49]. $\Delta$‘ denotes improvement.
@@
TABLE VI: Ablation study on our Object Jitter method. ‘COCO‘ denotes using the COCO [67] images for Object Jitter. ‘SA‘ denotes using 400k random images from the SA-1B [57] dataset for Object Jitter. ‘C.G.‘ denotes the CocoGlide dataset [49].
Num Object Jitter Operation Image Source CASIAv1 [46] Coverage [47] NIST16 [48] IMD20 [45] C.G. [49] CIMD [50] MISD [51] Average
@@
TABLE VII: Comparison study of our web-supervision on IML task. Blue number for improvement and Red number for decrease. ‘B.L.‘ denotes baseline setting. ‘PSCC-Synth.‘ denotes the synthetic dataset of PSCC-Net [56], ‘AIGC-Synth.‘ denotes the AIGC forgery dataset OpenSDID [21]. ‘O.J.‘ denotes our Object Jitter method.